OpenStudy (anonymous):

solve x2+3x+2=0 by using completing the square method

OpenStudy (anonymous):

x^2 + 2*3/2*x + 9/4 - 9/4 + 2 = 0 (x + 3/2)^2 = 9/4 - 2 => x + 3/2 = sqrt(1/4) x + 3/2 = |1/2| x + 3/2 = + 1/2 or x+ 3/2 = -1/2

jhonyy9 (jhonyy9):

- so i think x*x +x +2x +1*2 =0 x(x+1) +2(x+1) =0 (x+1)(x+2)=0